OpenStudy (anonymous):

you answer it.. and ill fan every last one of y'all but medal to whoever helps me.. Determine if the lines that pass through the given points are parallel, perpendicular or neither. Line A: (2, 3) and (--4, 5) Line B: (-4, 9) and (-3, 12)

OpenStudy (anonymous):

well do you know what those terms mean

OpenStudy (anonymous):

parallel, perpendicular

OpenStudy (anonymous):

GeoGebra.com If you use that it will give you a Graph and Plot everything for you :) Without needing a slope :)

OpenStudy (anonymous):

Find the slope of line A and the slope of line B. If both slopes are equal, the lines are parallel. I assume you can find the slope of a line given the two points. I will do A: slope = change of y/change of x = 5-3/-4-2 = 2/-6 = -1/3. Now you find the slope of line B.

OpenStudy (anonymous):

Also that Negative is that meant to be a Double Negative?

OpenStudy (anonymous):

well use the slope equation in order to find the slopes (y2-y1)/(x2-x1)

OpenStudy (anonymous):

then to find the equation of the line you can use point-slope form y-y1=m(x-x1) we can do this to make sure the lines are two separate lines

OpenStudy (anonymous):

remember parallel lines have the same slope but difference y-intercept and perpendicular lines have slopes that are negative reciprocals of each other

OpenStudy (anonymous):

Here :) I think this is what you are looking for :)